James has a 56% on his last math test. He got 14 correct. How many questions were on his test?
Answer by stanbon(75887)

James has a 56% on his last math test. He got 14 correct. How many questions were on his test?
----
Equation:
0.56*x = 14
x = 14/0.56 = 25 (# of test questions)
------------
Cheers,
Stan H.
--------------